# Scalable Gaussian Processes with Latent Kronecker Structure

### v1

@@ 2 Background and Related Work
where ${\mathbf{K}}_{{\mathbf{X}}{\mathbf{X}}}$ denotes the covariance matrix $[k({\mathbf{x}}_{i},{\mathbf{x}}_{j})]_{i,j=1}^{n}$ and ${\mathbf{K}}_{(\cdot){\mathbf{X}}}$, ${\mathbf{K}}_{{\mathbf{X}}(\cdot^{\prime})}$, and ${\mathbf{K}}_{(\cdot,\cdot^{\prime})}$ denote pairwise evaluations of the same kind. For a comprehensive introduction to GPs, we refer to Rasmussen & Williams (2006).
@@
While the above methods scale gracefully to large datasets, they are all fundamentally limited by performing inference of approximate GP models.
@@ 3 Latent Kronecker Structure
Since latent Kronecker structure can be understood as padding $n\leq pq$ observations to complete a grid of $pq$ values, a key question is at which point the number of missing values $pq-n$ is large enough such that this padding becomes inefficient and dense matrices without this structure are preferable.
@@ 4 Experiments
For this experiment, we consider the positions, velocities, and accelerations as $\mathcal{S}$ and the seven torques as $\mathcal{T}$. We choose $k_{\mathcal{S}}$ to be a squared exponential kernel and $k_{\mathcal{T}}$ to be a full-rank ICM kernel (Bonilla et al., 2007), demonstrating that LKGP is compatible with discrete kernels. We select subsets of the training data split with $p=5000$ joint positions, velocities, and accelerations, and their corresponding $q=7$ joint torques, and introduced $10$%, $20$%, …, $90$% missing values uniformly at random, resulting in $n\leq 35k$. The value of $p$ was chosen such that the kernel matrix without factorization fits into GPU memory. This results in a fairer comparison since it allows both methods to amortize kernel evaluation costs, such that the comparison is essentially between matrix multiplication costs (larger $p$ would force regular iterative methods to recompute kernel values, leading to even higher overall compute times). Details on the model fitting setup are provided in Appendix C.
Table 2: [Climate Data with Missing Values] Predictive performances and total runtimes of temperature and precipitation prediction across various missing data ratios. LKGP consistently achieves the best predictive performance and also requires the least time across both datasets and all missing ratios. Reported numbers are the mean over $5$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
For this experiment, we consider hyperparameter configurations as $\mathcal{S}$ and learning curve progression steps as $\mathcal{T}$, such that $p=2000$, $q=52$, and thus $n\leq 104000$ for each dataset. We choose both $k_{\mathcal{S}}$ and $k_{\mathcal{T}}$ to be squared exponential kernels. Out of the $p=2000$ curves per dataset, $10$% were provided as fully observed during training and the remaining $90$% were partially observed. The stopping point was chosen uniformly at random. See Figure 4 for an illustration and qualitative comparison of the GP methods which we considered for this experiment. Details on the model fitting setup are provided in Appendix C.
@@
For this experiment, we considered latitude and longitude as $\mathcal{S}$ and time in days as $\mathcal{T}$. We selected $p=5000$ spatial locations uniformly at random together with $q=1000$ days of temperature or precipitation measurements, starting on January 1st, 2020, resulting in two datasets of size $n\leq 5$M. See Figure 5 for an illustration. We choose $k_{\mathcal{S}}$ to be a squared exponential kernel and $k_{\mathcal{T}}$ to be the product of a squared exponential kernel and a periodic kernel to capture seasonal trends. Missing values were selected uniformly at random with ratios of $10$%, $20$%, …, $50$%. Details on the model fitting setup are provided in Appendix C.
@@ 5 Conclusion
The primary limitation of LKGP is that it employs a product kernel, which assumes that points are only highly correlated if they are close in both $\mathcal{S}$ and $\mathcal{T}$. While this is a reasonable assumption in many settings, such as correlated time series data, it will not always be appropriate. The other requirement for our approach to be applicable is that the data lives on at least a partial grid. This requirement is less restrictive since our model is highly competitive even if there are a lot of missing values, as our experiments and theory show. Furthermore, if the data does not even live on a partial grid, it would be possible to generate an artificial grid, for example via local interpolation.
@@ Appendix A Derivation of Asymptotic Break-even Points
Since $\gamma$ is the missing ratio, which is between 00 and $1$, we conclude that the asymptotic break-even point for time is
@@ Appendix B Additional Experiment Results
Table 3: Predictive performances and total runtimes of learning curve prediction on LCBench datasets $1-7$. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 4: Predictive performances and total runtimes of learning curve prediction on LCBench datasets $8-14$. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 5: Predictive performances and total runtimes of learning curve prediction on LCBench datasets $15-21$. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 6: Predictive performances and total runtimes of learning curve prediction on LCBench datasets $22-28$. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 7: Predictive performances and total runtimes of learning curve prediction on LCBench datasets $29-35$. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.

### v2

@@ 2 Background and Related Work
where ${\mathbf{K}}_{{\mathbf{X}}{\mathbf{X}}}$ denotes the covariance matrix $k({\mathbf{x}}_{i},{\mathbf{x}}_{j})]_{i,j=1}^{n}$ and ${\mathbf{K}}_{(\cdot){\mathbf{X}}}$, ${\mathbf{K}}_{{\mathbf{X}}(\cdot^{\prime})}$, and ${\mathbf{K}}_{(\cdot,\cdot^{\prime})}$ denote pairwise evaluations of the same kind. For a comprehensive introduction to GPs, we refer to Rasmussen & Williams (2006).
@@
While the above methods scale to large datasets, they are limited by performing inference of approximate GP models.
@@
State-Space Approaches
Hartikainen & Särkkä (2010); Sarkka et al. (2013) introduced state-space formulations for (spatio-)temporal GP regression, whose computation has a linear complexity with respect to the number of temporal observations for a given point in space. See Särkkä & Svensson (2023) for a review of the methods. Recently, Laplante et al. (2025) leveraged this formulation to accelerate robust spatiotemporal GP regression. Pförtner et al. (2025) extended the computation-aware framework of Wenger et al. (2024) to quantify errors caused by approximate filtering and smoothing methods. The main limitations of the state-space approaches to GP regression are 1) the requirement that the temporal kernel be stationary, and 2) that they scale quadratically or cubically with the spatial dimensionality.
The method proposed herein scales linearly in the spatial dimensionality, does not require the temporal kernel to be stationary, and is not limited to spatiotemporal data. If the method is applied to spatiotemporal data collected in uniform temporal intervals, and the temporal kernel is stationary, the method can be accelerated to be quasi-linear in the number of time steps by leveraging the Toeplitz structure of the temporal kernel matrix (Wilson & Nickisch, 2015).
@@ 3 Latent Kronecker Structure
Since latent Kronecker structure can be understood as padding $n\leq pq$ observations to complete a grid of $pq$ values, a key question is at which point the number of missing values pqnpq-n is large enough such that this padding becomes inefficient and dense matrices without this structure are preferable.
@@ 4 Experiments
For this experiment, we consider the positions, velocities, and accelerations as $\mathcal{S}$ and the seven torques as $\mathcal{T}$. We choose $k_{\mathcal{S}}$ to be a squared exponential kernel and $k_{\mathcal{T}}$ to be a full-rank ICM kernel (Bonilla et al., 2007), demonstrating that LKGP is compatible with discrete kernels. We select subsets of the training data split with $p=5000$ joint positions, velocities, and accelerations, and their corresponding $q=7$ joint torques, and introduced 1010%, 2020%, …, 9090% missing values uniformly at random, resulting in $n\leq 35k$. The value of $p$ was chosen such that the kernel matrix without factorization fits into GPU memory. This results in a fairer comparison since it allows both methods to amortize kernel evaluation costs, such that the comparison is essentially between matrix multiplication costs (larger $p$ would force regular iterative methods to recompute kernel values, leading to even higher overall compute times). Details on the model fitting setup are provided in Appendix C.
Table 2: [Climate Data with Missing Values] Predictive performances and total runtimes of temperature and precipitation prediction across various missing data ratios. LKGP consistently achieves the best predictive performance and also requires the least time across both datasets and all missing ratios. Reported numbers are the mean over 55 random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
For this experiment, we consider hyperparameter configurations as $\mathcal{S}$ and learning curve progression steps as $\mathcal{T}$, such that $p=2000$, $q=52$, and thus $n\leq 104000$ for each dataset. We choose both $k_{\mathcal{S}}$ and $k_{\mathcal{T}}$ to be squared exponential kernels. Out of the $p=2000$ curves per dataset, 1010% were provided as fully observed during training and the remaining 9090% were partially observed. The stopping point was chosen uniformly at random. See Figure 4 for an illustration and qualitative comparison of the GP methods which we considered for this experiment. Details on the model fitting setup are provided in Appendix C.
@@
For this experiment, we considered latitude and longitude as $\mathcal{S}$ and time in days as $\mathcal{T}$. We selected $p=5000$ spatial locations uniformly at random together with $q=1000$ days of temperature or precipitation measurements, starting on January 1st, 2020, resulting in two datasets of size $n\leq 5$M. See Figure 5 for an illustration. We choose $k_{\mathcal{S}}$ to be a squared exponential kernel and $k_{\mathcal{T}}$ to be the product of a squared exponential kernel and a periodic kernel to capture seasonal trends. Missing values were selected uniformly at random with ratios of 1010%, 2020%, …, 5050%. Details on the model fitting setup are provided in Appendix C.
@@ 5 Conclusion
The primary limitation of LKGP is that it employs a product kernel, which assumes that points are only correlated if they are close in both $\mathcal{S}$ and $\mathcal{T}$. This is a reasonable assumption in many settings, such as correlated time series data, but will not always be appropriate. The other requirement for our approach is that the data lives on a partial grid, which is less restrictive since our model is competitive even if there are a lot of missing values, as our experiments and theory show. Furthermore, if the data does not even live on a partial grid, it would be possible to generate an artificial grid, for example via local interpolation.
@@ Appendix A Derivation of Asymptotic Break-even Points
Since $\gamma$ is the missing ratio, which is between 0 and $1$, we conclude that the asymptotic break-even point for time is
@@ Appendix B Additional Experiment Results
Table 3: Predictive performances and total runtimes of learning curve prediction on LCBench datasets 171-7. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 4: Predictive performances and total runtimes of learning curve prediction on LCBench datasets 8148-14. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 5: Predictive performances and total runtimes of learning curve prediction on LCBench datasets 152115-21. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 6: Predictive performances and total runtimes of learning curve prediction on LCBench datasets 222822-28. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.
@@
Table 7: Predictive performances and total runtimes of learning curve prediction on LCBench datasets 293529-35. Reported numbers are the mean over $10$ random seeds. Results of the best model are boldfaced and results of the second best model are underlined.